1) What volume will a gas occupy at 740 mmpressure which at 1480 mm occupies 500cc?
2) At a given temperature the pressure of a gas reduces to 75% of its initial value and the volume increases by 40% of its initial value. Find temp. if the initial temp. was 10 degree celcius
1
Expert's answer
2013-08-06T08:07:02-0400
1) As a mathematical equation, Boyle's law is:
P1V1 = P2V2
where:
V - volume;
P - pressure
from this equation the volume will be:
V1 = P2V2/P1
V1 = 1480 mm * 500cc / 740 mm = 1000 cc;
Answer: 1000 cc

2) p1 = p1
v1= v1
t1=10c =283k
p2=(1-75/100)*p1
v2=(1-40/100)*v1
t2=?
p1*v1/t1=p2*v2/t2
t2=p2*v2*t1/p1*v1
t2=(1-75/100)*p1 * (1-40/100)*v1 * 283k/ p1*v1
from this:
t2=(1-75/100)* (1-40/100)* 283k;
t2 = -0,74*(-0,39) * 283k = 81.67 k
Answer: 81.67 k